OpenStudy (anonymous):

suppose the amount of radioactive substance left after t years is given by A(t)=50e^(-0.0125t).Find the half life of this radioactive substance

OpenStudy (anonymous):

The half-life is defined to be the amount of time needed for half the original amount to decay. The original amount is found when t = 0 (i.e. from when we start measuring). At t=0, A(0)=50...so this is your initial amount. At time t = u, half of the original amount will have decayed. So you'd have at this time, 25=50e^(-0.0125u) ---> 1/2=e^(-0.0125u) Take the natural log of both sides:\[\ln 1/2 = \ln (e^{-0.0125u}) \rightarrow -\ln 2 = -0.0125u \rightarrow u = \frac{\ln 2}{0.0125}\]i.e.\[u=8 \ln 2\]years. This is your half-life.

OpenStudy (anonymous):

so the half life is 8 ln 2 years?

OpenStudy (anonymous):

Yes. About 5.5yrs.

OpenStudy (anonymous):

thanks again